# Learn from Failure: Fine-Tuning LLMs with Trial-and-Error Data for Intuitionistic Propositional Logic Proving

### v1

@@ 1 Introduction
Figure 1: A simple example for how learning trial-and-error data impacts inference distribution.
@@
In this paper, we wish to empirically verify this intuition. To conduct the experiment, we would compare the conventional model trained on correct proof paths, and TrialMaster, the model trained on the whole proof tree, containing both correct paths and incorrect paths. See Figure  1(b). As such, TrialMaster can make predictions based on the failure information during inference time.
Since current open-source Automated Theorem Proving datasets do not contain complete proof trees, we create such a dataset, PropL, written in Lean. We focus on theorems of intuitionistic propositional logic. A simple example of an intuitionistic propositional logic theorem and its proof in Lean is shown below. The first line is the theorem statement, and the second and third lines are tactics that solve the theorem.
    theorem thm : p &$\rightarrow$\$& p := by
       intro h
       exact h
@@ 2 Related Work
Automated Theorem Proving. Automated Theorem Proving has evolved significantly since its inception, focusing mainly on developing computer programs that can autonomously prove mathematical theorems. Early ATP systems (mechanical theorem proving) were based on first-order logic Clocksin and Mellish (2003); Chang and Lee (2014), where the resolution method Robinson (1965) played a crucial role. Recent progress in Automated Theorem Proving has been marked by the integration of machine learning Bansal et al. (2019); Davies et al. (2021); Wagner (2021), especially LLMs Yang et al. (2023); Polu and Sutskever (2020); Han et al. (2021); Welleck et al. (2021); Jiang et al. (2022), and heuristic methods Holden and Korovin (2021), aimed at amplifying the efficiency and capacity of Automated Theorem Proving systems. Within the domain of LLMs, formal mathematical languages like Metamath Megill and Wheeler (2019), Lean de Moura et al. (2015), Isabelle Nipkow et al. (2002), and Coq Barras et al. (1997), serve as a bridge, enabling the precise expression and verification of mathematical theorems and concepts through a computer-verifiable format, thereby mitigating hallucination risks Nawaz et al. (2019). COPRA works on incorporating the backtracking information into the prompts and sends the prompt to GPT-4 without fine-tuning it to perform proof searching task Thakur et al. (2023). Baldur fine-tuned the LLMs with proofs and error information given by the proof assistant First et al. (2023). In contrast, our work focused on fine-tuning the LLMs with the complete past proof history without using the error message from the proof assistants.
@@
Propositional Logic Problem. Early implementations of ATP systems demonstrated the potential for computers to automate logical deductions, with notable examples including the Logic Theorist Crevier (1993); McCorduck (2004); Russell and Norvig (2010) and Gilmore’s program Davis (2001); Gilmore (1960). These systems laid the groundwork for the resolution of propositional logic problems, showcasing the ability of automated systems to handle logical reasoning tasks. Recent advancements in Automated Theorem Proving have revisited propositional logic problems, integrating modern computational techniques. Sekiyama et al. Sekiyama and Suenaga (2018) have employed Deep Neural Networks (DNNs) as a statistical approach to generate proofs for these theorems, while Kusumoto et al. Kusumoto et al. (2018) have explored graph representations coupled with reinforcement learning to find proofs. Furthermore, the sequence-to-sequence neural networks have been applied for deriving proof terms in intuitionistic propositional logic Sekiyama et al. (2017). This area of research is particularly intriguing due to the simplicity and importance of propositional logic in mathematics, and there is a growing interest in evaluating the capability of LLMs in tackling this specific mathematical domain.
@@
Trial-and-Error. The Chain-of-Thought (CoT) Wei et al. (2022); Wang et al. (2022); Zhou et al. (2022); Fu et al. (2022); Chu et al. (2023); Yu et al. (2023) approach, demonstrates that LLMs can be guided to perform step-by-step reasoning by incorporating intermediate reasoning steps in their prompts. This concept is expanded in later research, such as the Tree of Thoughts (ToT) Yao et al. (2023), which organizes reasoning into a tree structure, and the Graph of Thoughts (GoT) Besta et al. (2023), which adopts a graph format for thought structuring. Trial-and-error complements structured reasoning by allowing the model to empirically test hypotheses generated, thereby refining its reasoning process based on feedback from interactions or emulations. The Boosting of Thoughts (BoT) Anonymous (2024) prompting framework iteratively explores and evaluates multiple trees of thoughts to gain trial-and-error reasoning experiences, using error analysis from the LLMs to revise the prompt. We posit that teaching models to leverage both structured reasoning and trial-and-error methodologies can substantially improve their analytical and adaptive capabilities.
@@ 3.1 Data Generation of PropL
Theorem Generation. Consider the set of propositions $A$ with at most $p$ atomic propositions, $A$ can be inductively generated by the following grammar:
@@
where $P_{i}$ is the $i$-th atomic proposition with $1\leq i\leq p$. A connective $\land$, $\lor$, $\to$ is called an internal node of the proposition. We assign the following lexicographic order to propositions:
@@
For a fixed upper bound $p$ for the number of atomic propositions, we establish a bijection between the natural numbers and the set of propositional logic formulas. The counting can be made efficient using the Catalan Numbers Atkinson and Sack (1992). Figure 3 gives an example of mapping between propositions and natural numbers. Details about the encoding and decoding algorithms are provided in Appendix A.
@@
Proof Generation. Given a randomly sampled theorem, the proof of the theorem is constructed using the focusing method with polarization McLaughlin and Pfenning (2009); Liang and Miller (2009); Pfenning (2017). Proof search is divided into two stages: inversion and chaining. The inversion phase mechanically breaks down negative connectives (e.g. implications) in the goal and positive connectives (e.g. disjunctions) in the premises. After inversion, chaining will pick an implication in the premise or show one of the disjuncts in the conclusion, with backtracking. The proof search procedure terminates when the same atomic proposition appears in both the premise and the conclusion. An example of proof with trial-and-error information (backtracking) and that with trial-and-error information removed is shown in Figure 4.
@@
The polarization of the connectives affects the behavior of the inversion and the search procedure. We choose to uniformly polarize conjunctions that occur negatively (e.g. on the right-hand side of a sequence) as negative and polarize conjunctions that occur positively (e.g. on the left-hand side of an arrow) as positive. Atomic propositions are assigned polarities based on the connective that they first appear under.
@@
3.2 Construction of Training and Testing Sets
@@
We then split the 200,000 theorems into training and testing sets based on both of the word length distributions mentioned above. The word lengths of the proofs of the training data theorems fall within the lower 0.66 quantile of the two distributions of the word length of all the proofs of the 200,000 theorems (109,887 in total). The word lengths of the in-distribution testing also fall in that category (1000 in total). The word lengths of the proofs among the out-of-distribution testing theorems are above 0.8 quantile (1000 total) of the two distributions of the word lengths of all the proofs of the 200,000 theorems.
@@ 4.1 LLM Fine-Tuning
We utilize the training set formed in PropL for training both TrialMaster and the tactic generator in the DFS system. The numbers of theorems used in the training and testing datasets are presented in Table 1.
LLM fine-tuning with trial-and-error. In our approach, we randomly select two out of the shortest five among the ten proofs with trial-and-error information for each theorem in the training set and utilize them to train TrialMaster. Refer to Figure 1(b) describes this training process.
@@
Table 1: The scale of training and testing split in our PropL dataset.
@@
Training 109,887
In-dist. testing 1,000
Out-of-dist. testing 1,000
@@ 4.2 Inference
Inference method of model trained with trial-and-error. TrialMaster conducts inference on itself without any help from a backtracking system like DFS or BFS. It outputs two kinds of tactics: tactics in Lean and backtrack instruction. An example of a backtrack instruction would be like “no solution, return to state 2 [that leads to state 4]”, where state 4 is the current state. When TrialMaster is doing a proof search on the test set, it is prompted with all history paths, including previous tactics, states, the backtracking it made before, and the failed search path. It then outputs the entire proof path after. Nonetheless, we only utilize the first tactic in the output and employ Lean as a calculator to determine the next state, thereby ensuring the correctness of the state following the tactic. If the tactic output by TrialMaster is a backtrack instruction, it is then prompted with all the proof search history including the backtrack instruction and the state that the backtrack instruction instructs to return to. If that tactic is not a backtrack instruction, the tactic and the current state will be fed into Lean for producing the state after. TrialMaster is then prompted with the entire proof tree including the state that Lean calculated, and it should output a tactic again. This process is repeated until Lean identifies that the proof is complete or any Lean error occurs. We also note that TrialMaster only outputs one tactic at each state using greedy search.
@@ 5.1 Experiment setup
Hyperparameters. In the experiment, we vary temperature $t$ from $\{0.3,0.7,1.2,1.5,2.0\}$ and number of sampled tactics $N_{\text{sampled}}$ from $\{2,5,10,15,20\}$. We notice that in all experiments for temperature $t$ less than 1.5, there were only very few (less than 15) theorems that were still under search when the total steps for that search attempt reached 65. For $t$ higher than 1.5, $N_{\text{Lean}}$ starts to increase dramatically. At temperature 2 with $N_{\text{sampled}}$=20, $N_{\text{Lean}}$ climbs up to 32171 when the number of total steps reaches 65. Therefore, in our experiment, we set 65 as the search step limit to control time complexity.
@@
Figure 5: Experiment results on OOD task. (a) We fix $N_{\text{sampled}}=10$ to see the impact of temperature on the DFS system. (b) We fix $t=1.2$ to see the impact of the number of sampled tactics. The number of Lean calls is noted beside the marker. (c) Comparison of $N_{\text{Lean}}$ among our method and top 3 DFS systems with the highest success rate. In summary, training with trail-and-error achieves a higher success rate with a relatively lower search cost compared to the DFS systems.
@@ 5.2 Evaluation Metrics
Search cost. We define a metric to assess the search cost—the total number of Lean calls for tactic checking during proof searching for the entire testing set, denoted as $N_{\text{Lean}}$. This number truly reflected given the same proof search success rate, a lower $N_{\text{Lean}}$ indicates a more efficient system for proof searching. Note that backtracking instructions from our method do not require to apply Lean to check the state and, consequently do not add search cost.
@@ 5.3 Results and Analysis
TrialMaster outperforms conventional DFS system. We begin by evaluating the methods of the in-distribution testing set. Table 2 illustrates that both our method and the DFS system perform exceptionally well, achieving a success rate of nearly 100% in most configurations. This suggests that Llama-7b effectively masters in-distribution intuitionistic propositional logic theorems. Then, we compare the performance of the methods on the out-of-distribution task. The results are presented in Figure 5. Our method with trial-and-error significantly outperforms the DFS system across various hyperparameter configurations. Additionally, we observe that feeding more proofs without trial-and-error for LLM fine-tuning does not further improve the performance.
@@
TrialMaster achieves high success rates at lower search cost. For a direct comparison of search costs, we plot the $N_{\text{Lean}}$ values of our method alongside those of the top three DFS systems with the highest success rates among all the hyperparameters we experimented with, i.e., $t$=1.5, $N_{\text{sampled}}$=20 (87.2%), $t$=1.5, $N_{\text{sampled}}$ = 10 (86.1%), and $t$=1.8, $N_{\text{sampled}}$=10 (88.4%). This comparison is illustrated in Figure 4(c). Notably, we observe that the DFS systems that closely approach our model’s performance exhibit significantly higher search costs. With $t$=1.8, $N_{\text{sampled}}$=10, $N_{\text{Lean}}$ of the DFS system, which has 0.3% lower success rate than our method, $N_{\text{Lean}}$ has reached 31101, which is 72% higher than that of our method with trial-and-error. The high search cost makes the DFS system with high temperatures unfavorable. These results demonstrate that training with trial-and-error produces higher-quality tactics, achieving a higher success rate with relatively lower search cost.
Model learns backtracking capability from trial-and-error data. In the experiments, we find out that our TrialMaster successfully acquires the backtracking capability from proofs with trial-and-error information. This is evidenced by the fact that during TrialMaster’s proof search for theorems in the testing set, all backtracking instructions produced by the LLM adhere to the correct format and point to existing state numbers.

### v3

@@ 1 Introduction
Figure 1: A simple example for how learning trial-and-error data impacts inference distribution. See Figure  4 for a concrete case.
@@
In this paper, we wish to empirically verify this intuition. To conduct the experiment, we would compare the conventional model trained on correct proof paths, and TrialMaster, the model trained on the whole proof tree, containing both correct paths and incorrect paths. See Figure  2(b). As such, TrialMaster can make predictions based on the failure information during inference time.
Since current open-source Automated Theorem Proving datasets do not contain complete proof trees, we create such a dataset, PropL, written in Lean. We focus on theorems of intuitionistic propositional logic. A propositional logic formula in intuitionistic propositional logic $A$ is true if and only if $\vdash A$ has a derivation according to the rules of the intuitionistic propositional logic. These rules are listed in Figure 6 in Appendix A with explanations. This is to be distinguished from the classical logic, where a theorem of classical propositional logic admits a proof if and only if under all truth value assignments to all propositions that appear in the theorem, the theorem statement is evaluated to be true.
We give two elementary examples of theorems of intuitionistic propositional logic with proofs. We then give an example of a theorem whose proof contains a backtracking step. The first theorem admits a proof, where as the second doesn’t.
theorem_1 thm : p1 &$\rightarrow$\$& p1 $\vee$\$ p2 := by
intro h1
apply or.inl
exact h1
[…]
Our second example looks like the following.
theorem_2 thm : p1 &$\rightarrow$\$& p1 $\land$\$ p2 := by
The second theorem states that p1 implies p1 and p2. There is no proof to this theorem. When we assume p1, we cannot show both p1 and p2.
Proofs might include backtracking instructions. Here is another possible proof for theorem 1. After step 2, there is no possible proof step that can lead to final solution. Therefore we backtrack to step 1 and try a different proof step.
theorem_1 thm : p1 &$\rightarrow$\$& p1 $\vee$\$ p2 := by
intro h1 #this is step 1
apply or.inr #this is step 2
no solution, backtrack to step 1
apply or.inl
exact h1
Note that "no solution, backtrack to step 1" is not a tactic in Lean. It simply tells the system to backtrack to step 2 and start reasoning from there.
@@ 2 Related Work
Automated Theorem Proving. Automated Theorem Proving has evolved significantly since its inception, focusing mainly on developing computer programs that can autonomously prove mathematical theorems. Early ATP systems (mechanical theorem proving) were based on first-order logic Clocksin and Mellish (2003); Chang and Lee (2014), where the resolution method Robinson (1965) played a crucial role. Recent progress in Automated Theorem Proving has been marked by the integration of machine learning Bansal et al. (2019); Davies et al. (2021); Wagner (2021), especially LLMs Yang et al. (2023); Polu and Sutskever (2020); Han et al. (2021); Welleck et al. (2021); Jiang et al. (2022), and heuristic methods Holden and Korovin (2021), aimed at amplifying the efficiency and capacity of Automated Theorem Proving systems. Within the domain of LLMs, formal mathematical languages like Metamath Megill and Wheeler (2019), Lean de Moura et al. (2015), Isabelle Nipkow et al. (2002), and Coq Barras et al. (1997), serve as a bridge, enabling the precise expression and verification of mathematical theorems and concepts through a computer-verifiable format, thereby mitigating hallucination risks Nawaz et al. (2019). COPRA incorporates backtracking information into a prompt and sends the prompt to GPT-4 without fine-tuning it to perform proof search Thakur et al. (2023). Baldur fine-tunes LLMs with proofs and error information given by the proof assistant First et al. (2023). In contrast, our work focuses on fine-tuning LLMs with the complete past proof history without using the error message from the proof assistant.
@@
Propositional Logic Problem. Early implementations of ATP systems demonstrated the potential for computers to automate logical deductions, with notable examples including the Logic Theorist Crevier (1993); McCorduck (2004); Russell and Norvig (2010) and Gilmore’s program Davis (2001); Gilmore (1960). These systems laid the groundwork for the resolution of propositional logic problems, showcasing the ability of automated systems to handle logical reasoning tasks. Recent advancements in Automated Theorem Proving have revisited propositional logic problems, integrating modern computational techniques. Sekiyama et al. Sekiyama and Suenaga (2018) have employed Deep Neural Networks (DNNs) as a statistical approach to generate proofs for these theorems, while Kusumoto et al. Kusumoto et al. (2018) have explored graph representations coupled with reinforcement learning to find proofs. Furthermore, sequence-to-sequence neural networks have been applied for deriving proof terms in intuitionistic propositional logic Sekiyama et al. (2017). This area of research is particularly intriguing due to the simplicity and importance of propositional logic in mathematics, and there is a growing interest in evaluating the capability of LLMs in tackling this specific mathematical domain.
@@
Trial-and-Error. The Chain-of-Thought (CoT) Wei et al. (2022); Wang et al. (2022); Zhou et al. (2022); Fu et al. (2022); Chu et al. (2023); Yu et al. (2023) approach demonstrates that LLMs can be guided to perform step-by-step reasoning by incorporating intermediate reasoning steps in their prompts. This concept is expanded in later research, such as the Tree of Thoughts (ToT) Yao et al. (2023), which organizes reasoning into a tree structure, and the Graph of Thoughts (GoT) Besta et al. (2023), which adopts a graph format for thought structuring. Trial-and-error complements structured reasoning by allowing the model to empirically test hypotheses generated, thereby refining its reasoning process based on feedback from interactions or emulations. The Boosting of Thoughts (BoT) Chen et al. (2024) prompting framework iteratively explores and evaluates multiple trees of thoughts to gain trial-and-error reasoning experiences, using error analysis from the LLMs to revise the prompt. Lyra introduces correction mechanisms to improve the performance of the tactic generator Zheng et al. (2023).
@@ 3.1 Data Generation of PropL
Theorem Generation. Consider the set of propositions $A$ with at most $p$ atomic propositions. $A$ can be inductively generated by the following grammar:
@@
where $P_{i}$ is the $i$-th atomic proposition with $1\leq i\leq p$, $T$ stands for True, and $F$ stands for False. We note that an atomic proposition is just a statement like "Today is sunny". The concrete meaning of the statement is irrelevant with respect to the theorem proving task in this paper. A connective $\land$, $\lor$, $\to$ is called an internal node of the proposition. We assign the following lexicographic order to propositions:
@@
For a fixed upper bound $p$ for the number of atomic propositions, we establish a bijection between the natural numbers and the set of propositional logic formulas. The counting can be made efficient using the Catalan Numbers Atkinson and Sack (1992). Figure 3 gives an example of mapping between propositions and natural numbers. Details about the encoding and decoding algorithms are provided in Appendix B.
@@
Proof Generation. Given a randomly sampled theorem, the proof of the theorem is constructed using the focusing method with polarization McLaughlin and Pfenning (2009); Liang and Miller (2009); Pfenning (2017). Proof search is divided into two stages: inversion and chaining. The inversion phase mechanically breaks down negative connectives (e.g. implications) in the goal and positive connectives (e.g. disjunctions) in the premises. After inversion, chaining will pick an implication in the premise or show one of the disjuncts in the conclusion, with backtracking. The proof search procedure terminates when the same atomic proposition appears in both the premise and the conclusion. An example of proofs with trial-and-error information (backtracking) and with trial-and-error information removed are shown in Figure 4.
@@
The polarization of the connectives affects the behavior of the inversion and the search procedure. We choose to uniformly polarize conjunctions that occur negatively (e.g. on the right-hand side of an arrow) as negative and polarize conjunctions that occur positively (e.g. on the left-hand side of an arrow) as positive. Atomic propositions are assigned polarities based on the connective that they first appear under.
@@
3.2 Construction of Training and Test Sets
@@
We then split the 200,000 theorems into training and test sets based on both of the word length distributions mentioned above. The word lengths of the proofs of the training data theorems fall within the lower 0.66 quantile of the two distributions of the word length of all the proofs of the 200,000 theorems (109,887 in total). The word lengths of the in-distribution test set also fall in that category (1000 in total). The word lengths of the proofs among the out-of-distribution test theorems are above 0.8 quantile (1000 total) of the two distributions of the word lengths of all the proofs of the 200,000 theorems. We note that the baseline model and TrialMaster are trained on proofs of the same set of theorems.
@@ 4.1 LLM Fine-Tuning
We utilize the training set formed in PropL for training both TrialMaster and the tactic generator in the DFS system. The numbers of theorems used in the training and test datasets are presented in Table 1.
LLM fine-tuning with trial-and-error. In our approach, we randomly select two out of the shortest five among the ten proofs with trial-and-error information for each theorem in the training set and utilize them to train TrialMaster. Refer to Figure 2(b) for a description of this training process.
@@
Table 1: The scale of training and test split in our PropL dataset.
@@
Training set 109,887
In-dist. test set 1,000
Out-of-dist. test set 1,000
@@ 4.2 Inference
Inference method of model trained with trial-and-error. TrialMaster conducts inference on itself without any help from a backtracking system like DFS or BFS. It outputs two kinds of tactics: tactics in Lean and backtrack instructions. An example of a backtrack instruction would be like “no solution, return to state 2 [that leads to state 4]”, where state 4 is the current state. When TrialMaster is doing a proof search on the test set, it is prompted with all history paths, including previous tactics, states, the backtracking it made before, and the failed search path. It then outputs the entire proof path after. Nonetheless, we only utilize the first tactic in the output and employ Lean as a calculator to determine the next state, thereby ensuring the correctness of the state following the tactic. If the tactic outputted by TrialMaster is a backtrack instruction, it is then prompted with all the proof search history including the backtrack instruction and the state that the backtrack instruction says to return to. If that tactic is not a backtrack instruction, the tactic and the current state will be fed into Lean for producing the state after. TrialMaster is then prompted with the entire proof tree including the state that Lean calculated, and it should output a tactic again. This process is repeated until Lean identifies that the proof is complete or any Lean error occurs. We also note that TrialMaster only outputs one tactic at each state using greedy search.
@@ 5.1 Experimental Setup
Hyperparameters. In the experiment, we vary temperature $t$ from $\{0.3,0.7,1.2,1.5,2.0\}$ and number of sampled tactics $N_{\text{sampled}}$ from $\{2,5,10,15,20\}$. We notice that in all experiments for temperature $t$ less than 1.5, there were only very few (less than 15) theorems that were still under search when the total steps for that search attempt reached 65. For $t$ higher than 1.5, $N_{\text{Lean}}$ starts to increase dramatically. At temperature 2 with $N_{\text{sampled}}$=20, $N_{\text{Lean}}$ climbs up to 32,171 when the number of total steps reaches 65. Therefore, in our experiment, we set 65 as the search step limit to control time complexity.
@@
Figure 5: Experiment results on OOD task. (a) We fix $N_{\text{sampled}}=10$ to see the impact of temperature on the DFS system. (b) We fix $t=1.2$ to see the impact of the number of sampled tactics. The number of Lean calls is noted beside the marker. (c) Comparison of $N_{\text{Lean}}$ among our method and top 3 DFS systems with the highest success rate. In summary, training with trial-and-error achieves a higher success rate with a relatively lower search cost compared to the DFS systems.
@@ 5.2 Evaluation Metrics
Search cost. We define a metric to assess the search cost—the total number of Lean calls for tactic checking during proof search for the entire test set, denoted as $N_{\text{Lean}}$. Given the same proof search success rate, a lower $N_{\text{Lean}}$ indicates a more efficient system for proof search. Note that backtracking instructions from our method do not require applying Lean to check the state and, consequently do not add search cost.
@@ 5.3 Results and Analysis
TrialMaster outperforms conventional DFS system. We begin by evaluating the methods of the in-distribution test set. Table 2 illustrates that both our method and the DFS system perform exceptionally well, achieving a success rate of nearly 100% in most configurations. This suggests that Llama-7b effectively masters in-distribution intuitionistic propositional logic theorems. Then, we compare the performance of the methods on the out-of-distribution task. The results are presented in Figure 5. Our method with trial-and-error significantly outperforms the DFS system across various hyperparameter configurations. Additionally, we observe that feeding more proofs without trial-and-error for LLM fine-tuning does not further improve the performance.
@@
TrialMaster achieves high success rates at lower search cost. For a direct comparison of search costs, we plot the $N_{\text{Lean}}$ values of our method alongside those of the top three DFS systems with the highest success rates among all the hyperparameters we experimented with, i.e., $t$=1.5, $N_{\text{sampled}}$=20 (87.2%), $t$=1.5, $N_{\text{sampled}}$ = 10 (86.1%), and $t$=1.8, $N_{\text{sampled}}$=10 (88.4%). This comparison is illustrated in Figure 5(c). Notably, we observe that the DFS systems that closely approach our model’s performance exhibit significantly higher search costs. With $t$=1.8, $N_{\text{sampled}}$=10, $N_{\text{Lean}}$ of the DFS system, which has 0.3% lower success rate than our method, has reached 31,101, which is 72% higher than that of our method with trial-and-error. The high search cost makes the DFS system with high temperatures unfavorable. These results demonstrate that training with trial-and-error produces higher-quality tactics, achieving a higher success rate with relatively lower search cost.
Model learns backtracking capability from trial-and-error data. In the experiments, we find out that our TrialMaster successfully acquires the backtracking capability from proofs with trial-and-error information. This is evidenced by the fact that during TrialMaster’s proof search for theorems in the test set, all backtracking instructions produced by the LLM adhere to the correct format and point to existing state numbers.
Including failed search paths helps TrialMaster to learn. The following experiment shows that adding failed search paths to the training data for TrialMaster results in an overall gain. In this experiment, the model is only trained to learn the correct search paths and the backtracking instructions. The model is not trained to learn the failed search paths (we don’t compute the loss for the failed search paths during the training stage in this case). The proof success rate in this case is 75.6%, which is lower than TrialMaster’s proof success rate of 88.7%. The $N_{\text{Lean}}$ for the model is 13,600, which is lower than that of TrialMaster. This is expected since the model does not learn to predict failed search paths. Our explanation for why TrialMaster has a higher proof search success rate than the model trained in the previously mentioned experiment is that the failed search paths also contribute to improving the proof search success rate of the model. TrialMaster strategically tried some potentially failed search paths to gain a more comprehensive view of the problem, which then led to the correct search paths.